OpenStudy (anonymous):

Solve the equation: x/4 = 7/2

OpenStudy (anonymous):

The eqn x/4=7/2 can be written as 2x=28, thus x=14.

OpenStudy (anonymous):

May he power of christ compel you!

OpenStudy (anonymous):

O_O

OpenStudy (anonymous):

This Isnt Church...

OpenStudy (anonymous):

im talking about the user name of the asker!

OpenStudy (anonymous):

Lol God Needs To Help Him XD "666"

OpenStudy (anonymous):

BTW @MathHelper666 It's x=14

OpenStudy (anonymous):

Logic is absolutely right!

OpenStudy (anonymous):

Thanks!